OpenStudy (loser66):

Find inverse function of f(x) = x^2 (x in [1,4)]) please, help

OpenStudy (nincompoop):

y = x^2, solve for x then swap y and x

OpenStudy (loser66):

I need find : \[f:R\0\rightarrow R\\x\rightarrow \dfrac{1}{x^2}\] \[E=\{x \in R | 1\leq x\leq2\}\] find \(f^{-1}(E)\)

OpenStudy (loser66):

so that : let h(x) = 1/x--> h(E) =\(\{x\in R| 1\leq x\leq 2\}\) = [\(\dfrac{1}{2},1\)]

OpenStudy (nincompoop):

your R is suppose to be \( \sf \huge \mathbb{R} \)

OpenStudy (loser66):

and \(h^{-1}(x) = 1/x \) --> \(h^{-1}(E)=[1/2,1] \) also but \(f^{-1}(x) = (g(h(x))^{-1}= h^{-1}(g^{-1}(x) \)

OpenStudy (loser66):

Yes, that is Real

OpenStudy (loser66):

so that I need figure out what is \(g^{-1}(E)\)

OpenStudy (loser66):

first off, what is g^-(x)? is it \(g^{-1}(x) = \pm \sqrt{x}\)

OpenStudy (tkhunny):

On [1,4], it's just \(\sqrt{x}\). "Function" doesn't mean much when you deliberately write "\(\pm\)"

OpenStudy (nincompoop):

correct

OpenStudy (nincompoop):

it is important that you obtain the inverse function first, then test the conditions or range of values according to parameters

OpenStudy (loser66):

OK, thanks for the help. :)

OpenStudy (tkhunny):

Not quite. Rephrase. It is important that you obtain the inverse RELATION first, then test the conditions or range of values according to parameters SO THAT we have a function. There are two possible definitions for this f(x) on [1,4]. There is nothing sacred about choosing the positive one. Words mean things. Try to use them properly.